OpenStudy (anonymous):

Whats x and y if Y= x? and if Y=2x-4? PLEASE HELLPP

OpenStudy (anonymous):

Since y = x, you can replace the y in y = 2x -4 to make: x = 2x -4 ....which is an equation in one variable. Can you finish this off and solve for x?

OpenStudy (solomonzelman):

Y=2x-4 Y=2Y-4 -Y=4 Y=4 since x=y therefore x=4 just like y=4

OpenStudy (anonymous):

thank you. i just moved schools and missed a chunk of math which was slope everything so iv been pretty confused

OpenStudy (solomonzelman):

Do you understand what I did?

OpenStudy (anonymous):

Yes i got it! Now i need to graph it.. cause its solve each system of equations by graphing

OpenStudy (solomonzelman):

well you know that x=y x=y looks like this 

OpenStudy (solomonzelman):

where (x,y) and x=y so the coordinates can be (-111,-111) (2,2) (4,4) -7,-7) as long as x=y you get the idea...